# Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination

### v2

@@ 4 Synthetic Dataset
We created a novel synthetic dataset of indoor scenes (see Fig. 5) to train our network, which is based on the 3D-Front [11] public interior dataset with over 60k distinct house layouts. To the best of our knowledge, there is no publicly available dataset containing completely annotated lighting (with factorized shading and reflectance, separated direct/indirect and diffuse/specular components) and geometry features that adapt well to our task. Our dataset will be released publicly online upon publication to fill the absence of a high-quality dataset for learning-based shading prediction tasks.
Our synthetic dataset contains over 30,000 high dynamic range (HDR) images rendered from 1.5k indoor layouts from 3D-Front, each with pixel-wise annotations of lighting, texture, and geometry features. While all textures in 3D-Front are restricted to diffuse materials, and some surfaces have no materials, we also assign physically based materials to the surfaces, which provide highly diverse visual effects and improve visual realism compared to similar synthetic datasets as [39, 50]. Note that our dataset also applies to many other computer vision tasks, such as indoor scene understanding, inverse rendering, intrinsic decomposition, etc. Please see the supplementary document for more details and illustrations about material assignment and dataset generation.
To create photo-realistic rendered images with less manual effort, we use a custom data generation pipeline based on Blender Cycles [4], which can be summarized as the following steps:
Texture assignment To achieve more realism on our synthetic dataset, we use an open-source CC0 texture library [8] which includes over 1,300 types of physically based materials covering various surfaces such as wood, metal, glass, etc. We randomly assign certain materials based on the object type in a rule-based style, using a similar strategy with [9]. For example, for each wall surface, we uniformly sample one material among all wood, marble, or concrete-type materials. This rule-based technique for random texture assignment significantly reduces the effort of manually assigning the textures for each object while preserving sufficient consistency and diversity for indoor layouts.
@@ 5.1 Implementation Detail
We implement our network model using PyTorch. For the output of each convolutional layer, we apply Leaky ReLU activation, except for the last layer where an exponential activation is used. For both the generator and discriminator, we use Adam optimizer [23] for optimization and Xavier [12] for network initialization. Dropout [40] is used in up-sampling layers to avoid over-fitting. The model is trained for 80 epochs using over 50 hours. Model performance is measured on an RTX 3090, using $768\times 512$ (3:2 ratio) as the standard resolution for network inference to synthesize image. Please refer to the supplemental material for our code implementation and the pre-trained model, which will be released online along with the training dataset.
@@ 5.2 Results and Discussions
Pix2pix [18] 41.83M 0.9611 28.88 0.0352
Nalbach et al. [32] 7.75M 0.9564 27.95 0.0391
Xin et al. [46] 3.95M 0.9658 30.28 0.0312
@@
We use three evaluation metrics to compare the performance of different methods, including SSIM, PSNR, and LPIPS [49]. The LPIPS metric calculates perceptual similarity based on high-level features in a pre-trained AlexNet [25], which is more robust to noise and better resembles human choice. We compare two recent works on global illumination prediction, including Nalbach et al. [32] and Xin et al. [46]. Our method adopts a neural rendering approach based on shading space, predicting indirect illumination from screen space to achieve real-time HDR global illumination estimation. Its core idea is to efficiently simulate long-range indirect light transport through an image-to-image learning paradigm, without relying on explicit geometric modeling or physical path integration, thus belonging to a fast, screen-space neural rendering framework. In contrast, [51] focuses on 3D explicit scene modeling and object-level light transport representation. It implicitly models the illumination contribution of each object in the scene, superposing them in latent space to solve for global illumination. These methods differ fundamentally in technical routes, input spaces, and problem formulations. Therefore, such scene-space neural rendering methods [51, 41, 14] are not included for direct comparisons.
@@
Note that Nalbach et al. [32] directly reconstructs the global illumination image, while Xin et al. [46] predicts the indirect component, similar to our task setup. For Nalbach et al.’s implementation, we use the "mono" version they described. We train all models on our dataset since the dataset they used is not publicly available. We also compare our method with the commonly used models in image-to-image translation or image style transfer tasks, including Pix2pix [18] and ResNet [15] with the same training setup and loss function as our model. All evaluation results are obtained on our test set, including 1000 rendered frames from our synthetic dataset and about 200 frames from public scenes we collected, e.g., PinkRoom, LivingRoom, Gallery and Conference. The averaged results of the evaluation metrics are shown in Tab. 1.
@@
As can be seen, our model consistently outperforms all previous works and different network models on the three metrics. While the SSIM metric has a somewhat smaller difference among different methods, there is a huge gap between our method and other methods on the perceptual metric LPIPS, implying more natural illumination produced by our method. Visualizations of the whole predicted illumination are shown in Fig. 6. We further look into the predicted indirect component (demodulated from texture) in Fig. 7, which is more intuitive for comparing the performance of different methods. As observed, Xin et al.’s method used a lightweight model and tended to produce over-smoothed illumination on high-frequency regions, while the deeper models (Pix2pix and ResNet) using more parameters sometimes produced noisy and distorted results. Our method strikes a balance between model complexity and image quality, producing more plausible results at a reasonable time budget ($12\text{ms}$ at $768\times 512$), compared to the methods of Nalbach et al. [32] ($6\mathrm{ms}$) and Xin et al. [46] ($13\mathrm{ms}$).
@@ 5.3 Validation
Effectiveness of monochromatic design We regard the monochromatic design as the main reason that helps the model capture the correct color from indirectly reflected light (see Fig. 11). To further validate this, we implement an alternate version of our model, taking in the RGB inputs and directly reconstructing the RGB output, namely the Normal network. Both versions of the network use the same architecture except for the image reconstruction process. As shown in Fig. 10, our model is significantly more capable of capturing the colored indirect lighting. This is most visible in the regions consisting of indirect illumination only (e.g., the shadowed areas in CornellBox). Note that while Nalbach et al. [32] also suggested the independent inference for each channel, they simply take it in a completely independent manner, resulting in visual artifacts like color-shifting (see Fig. 8 in the main text and also Fig. 5 in Nalbach et al. [32]) due to their network’s poor ability to preserve linearity (i.e., $\mathbf{Net}(\alpha x)\neq\alpha\mathbf{Net}(x)$). In contrast, our network design takes several strategies to avoid this downside, including reusing shared geometry features, applying an inter-channel perceptual loss, and an effective data augmentation strategy for HDR training (see Sec. 3).
@@ 5.4 Discussion
GPU-accelerated ray tracing alternatives While there are multiple existing methodologies for efficient specular or glossy reflections in different use cases [29, 17], the diffuse indirect illumination needs more sophisticated processing due to its uniform scattering nature (resulting by its hemispherical diffuse BRDF lobe, instead of glossy or Dirac delta ones). Recently, the development of GPU ray tracing acceleration offered new capability for real-time indirect illumination, among which sparsely traced results with denoising is a promising technique.
@@
Figure 14: Our method vs. denoised path-tracer (PT). Path tracing ($1\mathrm{spp}$) with next event estimation (NEE) is used, which is then denoised with the AI-Accelerated denoiser in OptiX [3]. Our method is closer to the reference in many scenarios in terms of the predicted indirect component while using less time consumption than the Denoised PT.
Rendering Pipeline Integration Our model can be integrated into an interactive rendering pipeline, given that the direct illumination is relatively simple and can be approximated with real-time algorithms [10, 16]. By doing this, all the inputs are generated on-the-fly within GPU memory, which can be immediately sent to the neural networks without the additional cost of memory transfer operations. However, the practicality of our method is limited in the following aspects, which could be improved straightforwardly:
@@ 6 Conclusion, Limitation, and Future Works
In the future, we will investigate more effective techniques for indirect illumination prediction while maintaining a relatively lightweight and efficient network architecture for VR applications, especially binocular rendering. Meanwhile, better exploiting 3D spatial features or implicit scene representation (e.g., via multi-view images or temporal coherence) may also be helpful for illumination prediction. Our proposed approaches are also potentially beneficial for other tasks. Specifically, the geometry-aware aggregation module might be preferred when geometry-supervised long-range modeling is needed. Moreover, while the monochromatic design is hardly seen in computer vision tasks, it would be interesting to see if it could be applied to various neural rendering tasks, e.g., learning-based Monte Carlo denoising [3] or neural radiance fields [30].
Acknowledgments
This work was supported by Southern Marine Science and Engineering Guangdong Laboratory (Zhuhai) (No. SML2021SP101) and Shenzhen Science and Technology Program (KJZD20230923114114028).

### v3

@@ 4 Synthetic Dataset
We created a novel synthetic dataset of indoor scenes (see Fig. 5) to train our network, which is based on the 3D-Front [fu20203dfront] public interior dataset with over 60k distinct house layouts. To the best of our knowledge, there is no publicly available dataset containing completely annotated lighting (with factorized shading and reflectance, separated direct/indirect and diffuse/specular components) and geometry features that adapt well to our task. Our dataset will be released publicly online upon publication to fill the absence of a high-quality dataset for learning-based shading prediction tasks.
Our synthetic dataset contains over 30,000 high dynamic range (HDR) images rendered from 1.5k indoor layouts from 3D-Front, each with pixel-wise annotations of lighting, texture, and geometry features. While all textures in 3D-Front are restricted to diffuse materials, and some surfaces have no materials, we also assign physically based materials to the surfaces, which provide highly diverse visual effects and improve visual realism compared to similar synthetic datasets as [song2017suncg, zhang2017pbrs]. Note that our dataset also applies to many other computer vision tasks, such as indoor scene understanding, inverse rendering, intrinsic decomposition, etc. Please see the supplementary document for more details and illustrations about material assignment and dataset generation.
To create photo-realistic rendered images with less manual effort, we use a custom data generation pipeline based on Blender Cycles [blender], which can be summarized as the following steps:
Texture assignment To achieve more realism on our synthetic dataset, we use an open-source CC0 texture library [cc0textures] which includes over 1,300 types of physically based materials covering various surfaces such as wood, metal, glass, etc. We randomly assign certain materials based on the object type in a rule-based style, using a similar strategy with [denninger2019blenderproc]. For example, for each wall surface, we uniformly sample one material among all wood, marble, or concrete-type materials. This rule-based technique for random texture assignment significantly reduces the effort of manually assigning the textures for each object while preserving sufficient consistency and diversity for indoor layouts.
@@ 5.1 Implementation Detail
We implement our network model using PyTorch. For the output of each convolutional layer, we apply Leaky ReLU activation, except for the last layer where an exponential activation is used. For both the generator and discriminator, we use Adam optimizer [kingma2014adam] for optimization and Xavier [glorot2010xavier] for network initialization. Dropout [srivastava2014dropout] is used in up-sampling layers to avoid over-fitting. The model is trained for 80 epochs using over 50 hours. Model performance is measured on an RTX 3090, using $768\times 512$ (3:2 ratio) as the standard resolution for network inference to synthesize image. Please refer to the supplemental material for our code implementation and the pre-trained model, which will be released online along with the training dataset.
@@ 5.2 Results and Discussions
Pix2pix [isola2017image] 41.83M 0.9611 28.88 0.0352
Nalbach et al. [nalbach2017deep] 7.75M 0.9564 27.95 0.0391
Xin et al. [xin2021lightweight] 3.95M 0.9658 30.28 0.0312
@@
We use three evaluation metrics to compare the performance of different methods, including SSIM, PSNR, and LPIPS [zhang2018lpips]. The LPIPS metric calculates perceptual similarity based on high-level features in a pre-trained AlexNet [krizhevsky2012alexnet], which is more robust to noise and better resembles human choice. We compare two recent works on global illumination prediction, including Nalbach et al. [nalbach2017deep] and Xin et al. [xin2021lightweight]. Our method adopts a neural rendering approach based on shading space, predicting indirect illumination from screen space to achieve real-time HDR global illumination estimation. Its core idea is to efficiently simulate long-range indirect light transport through an image-to-image learning paradigm, without relying on explicit geometric modeling or physical path integration, thus belonging to a fast, screen-space neural rendering framework. In contrast, [zheng2024neural] focuses on 3D explicit scene modeling and object-level light transport representation. It implicitly models the illumination contribution of each object in the scene, superposing them in latent space to solve for global illumination. These methods differ fundamentally in technical routes, input spaces, and problem formulations. Therefore, such scene-space neural rendering methods [zheng2024neural, su2024dnr, su2025vfnr, hadadan2021neural] are not included for direct comparisons.
@@
Note that Nalbach et al. [nalbach2017deep] directly reconstructs the global illumination image, while Xin et al. [xin2021lightweight] predicts the indirect component, similar to our task setup. For Nalbach et al.’s implementation, we use the "mono" version they described. We train all models on our dataset since the dataset they used is not publicly available. We also compare our method with the commonly used models in image-to-image translation or image style transfer tasks, including Pix2pix [isola2017image] and ResNet [he2016deep] with the same training setup and loss function as our model. All evaluation results are obtained on our test set, including 1000 rendered frames from our synthetic dataset and about 200 frames from public scenes we collected, e.g., PinkRoom, LivingRoom, Gallery and Conference. The averaged results of the evaluation metrics are shown in Tab. 1.
@@
As can be seen, our model consistently outperforms all previous works and different network models on the three metrics. While the SSIM metric has a somewhat smaller difference among different methods, there is a huge gap between our method and other methods on the perceptual metric LPIPS, implying more natural illumination produced by our method. Visualizations of the whole predicted illumination are shown in Fig. 6. We further look into the predicted indirect component (demodulated from texture) in Fig. 7, which is more intuitive for comparing the performance of different methods. As observed, Xin et al.’s method used a lightweight model and tended to produce over-smoothed illumination on high-frequency regions, while the deeper models (Pix2pix and ResNet) using more parameters sometimes produced noisy and distorted results. Our method strikes a balance between model complexity and image quality, producing more plausible results at a reasonable time budget ($12\text{ms}$ at $768\times 512$), compared to the methods of Nalbach et al. [nalbach2017deep] ($6\mathrm{ms}$) and Xin et al. [xin2021lightweight] ($13\mathrm{ms}$).
@@ 5.3 Validation
Effectiveness of monochromatic design We regard the monochromatic design as the main reason that helps the model capture the correct color from indirectly reflected light (see Fig. 11). To further validate this, we implement an alternate version of our model, taking in the RGB inputs and directly reconstructing the RGB output, namely the Normal network. Both versions of the network use the same architecture except for the image reconstruction process. As shown in Fig. 10, our model is significantly more capable of capturing the colored indirect lighting. This is most visible in the regions consisting of indirect illumination only (e.g., the shadowed areas in CornellBox). Note that while Nalbach et al. [nalbach2017deep] also suggested the independent inference for each channel, they simply take it in a completely independent manner, resulting in visual artifacts like color-shifting (see Fig. 8 in the main text and also Fig. 5 in Nalbach et al. [nalbach2017deep]) due to their network’s poor ability to preserve linearity (i.e., $\mathbf{Net}(\alpha x)\neq\alpha\mathbf{Net}(x)$). In contrast, our network design takes several strategies to avoid this downside, including reusing shared geometry features, applying an inter-channel perceptual loss, and an effective data augmentation strategy for HDR training (see Sec. 3).
@@ 5.4 Discussion
GPU-accelerated ray tracing alternatives While there are multiple existing methodologies for efficient specular or glossy reflections in different use cases [mcguire2014efficient, hirvonen2019accurate], the diffuse indirect illumination needs more sophisticated processing due to its uniform scattering nature (resulting by its hemispherical diffuse BRDF lobe, instead of glossy or Dirac delta ones). Recently, the development of GPU ray tracing acceleration offered new capability for real-time indirect illumination, among which sparsely traced results with denoising is a promising technique.
@@
Figure 14: Our method vs. denoised path-tracer (PT). Path tracing ($1\mathrm{spp}$) with next event estimation (NEE) is used, which is then denoised with the AI-Accelerated denoiser in OptiX [chaitanya2017mc]. Our method is closer to the reference in many scenarios in terms of the predicted indirect component while using less time consumption than the Denoised PT.
Rendering Pipeline Integration Our model can be integrated into an interactive rendering pipeline, given that the direct illumination is relatively simple and can be approximated with real-time algorithms [fernando2005pcss, heitz2016real]. By doing this, all the inputs are generated on-the-fly within GPU memory, which can be immediately sent to the neural networks without the additional cost of memory transfer operations. However, the practicality of our method is limited in the following aspects, which could be improved straightforwardly:
@@ 6 Conclusion, Limitation, and Future Works
In the future, we will investigate more effective techniques for indirect illumination prediction while maintaining a relatively lightweight and efficient network architecture for VR applications, especially binocular rendering. Meanwhile, better exploiting 3D spatial features or implicit scene representation (e.g., via multi-view images or temporal coherence) may also be helpful for illumination prediction. Our proposed approaches are also potentially beneficial for other tasks. Specifically, the geometry-aware aggregation module might be preferred when geometry-supervised long-range modeling is needed. Moreover, while the monochromatic design is hardly seen in computer vision tasks, it would be interesting to see if it could be applied to various neural rendering tasks, e.g., learning-based Monte Carlo denoising [chaitanya2017mc] or neural radiance fields [mildenhall2020nerf].